# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v1

@@ VI-B Conditions of equality between efficient SFIM and parametric efficient FIM
By collecting the previous results, we are ready to state the following proposition whose proof is given in Appendix 203:
@@ VI-D The elliptical parameterized “low-rank” model
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\mathbf{0},\bm{\Sigma}_{0},g_{0})$ be a zero-mean RES-distributed vector whose scatter matrix $\bm{\Sigma}_{0}$ is modeled as:
@@
  • $\mathbf{A}_{0}\in\mathbb{R}^{m\times p}$ is a full rank column matrix with $m>p$, where $\bm{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{q}$ collects the parameter of interest and characterizes $\mathbf{A}_{0}$ and where the function $\mathbf{A}(\bm{\gamma})$ is differentiable,
@@ VII-A Real-complex representations
$p_{X}(\mathbf{x}|\bm{\mu},\bm{\Sigma},{\bf\Omega},g_{c})=|\widetilde{\bf\Sigma}|^{-1/2}g_{c}\left[2^{-1}(\widetilde{\bf x}-\widetilde{\bm{\mu}})^{H}\widetilde{\bf\Sigma}^{-1}(\widetilde{\bf x}-\widetilde{\bm{\mu}})\right],$ (129)
@@
    Use the complex Hilbert space $(\mathcal{H},\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}})$ in (1), the “complex-aware” inner product $\left\langle h_{1},h_{2}\right\rangle_{\mathcal{H}}\triangleq{E_{0}\{h_{1}h_{2}^{*}\}}$ to express projection operators and tangent spaces,
@@
for NC-CES distributions where $\beta(g_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}\varphi^{2}_{c,0}(\mathcal{Q}_{c})\}}{m}$ with $\varphi_{c,0}(t)\triangleq\frac{-1}{g_{c,0}(t)}\frac{dg_{c,0}(t)}{dt}$ and $[\mathbf{J}[\widetilde{\bm{\mu}}_{0}]]_{i,j}\triangleq\left.\frac{\partial[\widetilde{\bm{\mu}}(\bm{\theta})]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}}$. In the particular case of C-CES distributions, it is immediate to verify that the FIM in (134) reduces to
@@
$[{\bf J}[\mathrm{vec}({\bf A}_{0})]]_{i,j}\triangleq\left.\frac{\partial[\mathrm{vec}({\bf A}(\bm{\gamma}))]_{i}}{\partial\gamma_{j}}\right|_{\bm{\gamma}=\bm{\gamma}_{0}}$ and where here $\alpha(g_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}^{2}\varphi^{2}_{c,0}(\mathcal{Q}_{c})\}}{m(m+1)}$. Note that the parametric efficient FIM (137) reduces to:
@@ VIII Conclusion
    $\begin{split}||\mathbf{h}-\mathbf{u}||_{\mathcal{H}_{q}}&=\sqrt{\left\langle\mathbf{h}-\mathbf{u},\mathbf{h}-\mathbf{u}\right\rangle}_{\mathcal{H}_{q}}\\ &=\sqrt{\sum\nolimits_{i=1}^{q}\left\langle h_{i}-u_{i},h_{i}-u_{i}\right\rangle_{\mathcal{H}}}=\sqrt{\sum\nolimits_{i=1}^{q}||h_{i}-u_{i}||_{\mathcal{H}}}.\end{split}$ (147)
    As a consequence, minimizing $||\mathbf{h}-\mathbf{u}||_{\mathcal{H}_{q}}$ is equivalent to minimize each term $||h_{i}-u_{i}||_{\mathcal{H}}$. Then, the equality (145) follows from the definition of orthogonal projection in $\mathcal{H}$ onto $\mathcal{U}$. ∎
@@
    where $[\mathbf{G}(\mathbf{a},\mathbf{b})]_{i,j}\triangleq\left\langle a_{i},b_{j}\right\rangle_{\mathcal{H}}$.
@@
    To show that $[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}$ is orthogonal to $p_{j}$, $\forall i,j$, we note that, according to its definition given in (18), $[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}$ is the residual of $[\mathbf{s}_{\bm{\gamma}_{0}}]_{i}$ after projection onto $\mathcal{T}_{2}+\mathcal{T}_{3}$. As a direct consequence, we have that $[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}\perp(\mathcal{T}_{2}+\mathcal{T}_{3})$, or equivalently, $[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}\in(\mathcal{T}_{2}+\mathcal{T}_{3})^{\perp}$. Moreover, again by definition in (25), $p_{j}\in(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}\subseteq\mathcal{T}_{2}+\mathcal{T}_{3}$. Consequently, $\left\langle[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i},p_{j}\right\rangle=0$, $\forall i,j$ since $(\mathcal{T}_{2}+\mathcal{T}_{3})\perp(\mathcal{T}_{2}+\mathcal{T}_{3})^{\perp}$.
@@
    By its definition, given in (13), $[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$. As a consequence, we have that the first projection in the RHS of (154) can be evaluated as $\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}^{\perp})=\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$. And therefore $\mathbf{p}=\bm{0}\Leftrightarrow\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp})\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$ and $[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{3}^{\perp}$. Then because $[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$, $\mathbf{p}=\bm{0}\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{3}^{\perp}$ or equivalently, $[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\perp\mathcal{T}_{3}$.
@@
    $\left(\frac{d[\mathbf{V}_{S}]_{11}}{d\underline{\mathrm{vecs}}(\mathbf{V}_{S})}\right)^{T}\triangleq\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}=-\frac{\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m})_{1}}.$ (160)
@@
    $\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}=-\frac{\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m})_{1}}.$ (162)
@@
    We note that, for $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$, we have $\mathbf{D}_{S}^{\bm{\Sigma}}=\mathbf{e}_{1}\mathbf{e}_{1}^{T}$, $\mathbf{D}_{S}^{\bm{\Sigma}}=m^{-1}\mathbf{I}_{m}$ and $\mathbf{D}_{S}^{\bm{\Sigma}}=m^{-1}|\bm{\Sigma}|^{1/m}\bm{\Sigma}^{-1}$ respectively.
@@
      $[\mathbf{M}_{S}^{\bm{\Sigma}}]^{T}\underline{\mathrm{vecs}}(\mathbf{A})=\mathbf{D}_{m}\mathbf{K}_{\mathbf{V}_{S,0}}\underline{\mathrm{vecs}}(\mathbf{A})=\mathbf{D}_{m}\mathrm{vecs}(\mathbf{A})=\mathrm{vec}(\mathbf{A}).$ (164)
@@
      $\mathbf{M}_{S}^{\bm{\Sigma}}$ has full row rank equal to $m(m+1)/2-1$.
@@
    The proof of these properties can be found in [10, Lemma 4.2]. Let us now focus on the property P4 to add some insight on the image of $\mathbf{M}_{S}^{\bm{\Sigma}}$. From the Property P3, we have that the dimension of $\text{Im}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$, i.e. the number of the linearly independent rows of $\mathbf{M}_{S}^{\bm{\Sigma}}$ is equal to $m(m+1)/2-1$. Moreover, as a direct consequence of the rank-nullity theorem, we have the following additional property:
@@
      An orthonormal basis of the $m(m+1)/2-1$-dimensional image $\text{Im}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ of the restriction to $\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ of the linear application defined by $\mathbf{M}_{S}^{\bm{\Sigma}}$ is given by the columns of the matrix $\mathbf{V}_{\bm{\Sigma}}=\mathbf{D}_{m}\mathbf{U}_{\bm{\Sigma}}\in\mathbb{R}^{m^{2}\times(m(m+1)/2-1)}$ such that :
@@
[…]
@@
    where $\mathbf{s}_{\bm{\rho}_{0},i}=\nabla_{\bm{\rho}}\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho}_{0},i})$ is the score vector of $\bm{\rho}_{0}\in\Upsilon_{i}$ in the $i-th$ parametric sub-model $\mathcal{P}_{\tau_{i}}$. Consequently, according to [14] and [13, Sec. 4.4], the tangent space $\mathcal{T}_{g_{0}}$ can be defined as the closure  [^4] The closure $\overline{\mathcal{A}}$ of a set $\mathcal{A}$ is defined as the smallest closed set that contains $\mathcal{A}$, or equivalently, as the set of all elements in $\mathcal{A}$ together with all the limit points of $\mathcal{A}$. of the union of all the (parametric) tangent spaces $\mathcal{P}_{\tau_{i}}$:
@@
    Equivalently, $\mathcal{T}_{g_{0}}\subseteq\mathcal{H}$ is the subspace  [^5] The closure of a union of linear spaces doesn’t need to be linear, in general. However, as discussed in [50, Assumption S] and [13, Sec. 4.4, Remark 5]), $\mathcal{T}_{g_{0}}$ is a linear space in the vast majority of the non-pathological statistical models. of $\mathcal{H}$ composed by all the functions $h\in\mathcal{T}_{g_{0}}\subseteq\mathcal{H}$ for which there exists a sequence $\{\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i},\mathbf{c}_{i}\in\mathbb{R}^{r_{i}}\}_{i\in\mathbb{N}}$ such that $||h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}}||^{2}=E_{0}\left\{(h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}})^{2}\right\}\rightarrow 0$.
@@
    Consequently, $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}$ is clearly an element of $\mathcal{T}_{g_{0}}$ thanks to the linearity of the expectation operator and to the fact that $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\}=0$ and $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]^{2}_{j}\}<+\infty$, for $j=1,\ldots,r_{i}$ from usual properties of the score vectors (under the regularity conditions discussed in [17, Sects. 6.2, 6.3]). Then $[\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\in\mathcal{H}$ for $j=1,\ldots,r_{i}$ and it is measurable w.r.t. $Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})$, then $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}\in\mathcal{T}_{g_{0}},\forall i\in\mathbb{N}$.
@@
[…]
@@
    Consequently, since, as said before, any element of $\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$ is of the form $\mathbf{c}^{T}\mathbf{s}_{\bm{\rho}_{0},\tilde{i}}$ for some $\mathbf{c}\in\mathbb{R}^{r_{\tilde{i}}}$, we just need to choose $\mathbf{c}=\mathbf{e}_{j}$ to prove that $\mathcal{T}_{g_{0}}\ni\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\in\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$, where $\tilde{h}_{j}$ is a bounded function. The proof is completed by noticing that the set of bounded functions is dense in $\mathcal{H}$ and consequently, any element $h\in\mathcal{T}_{g_{0}}$ can be obtained as a converging sequence of bounded functions. This allow us to state that $\mathcal{T}_{g_{0}}\subset\overline{\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}}$ and conclude the proof. ∎
@@
    $[\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]^{-1}=2\alpha^{-1}(g_{0})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{2\alpha^{-1}(g_{0})(\alpha(g_{0})-1)}{(m+2)\alpha(g_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}.$ (184)
@@
    Proof of Proposition VI.3. The different steps of the proof are based on some reasoning and notations of [44, Sec. VI of supplement material] for C-CES distributions and the general ”low-rank” model (137), that itself takes up the steps of the proof presented in [51].
@@
    which can be written, by means of the notation used in [44, Sec; V of supplement material], in the following compact form
@@
    with ${\bf u}_{n}={\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\mathrm{vec}({\bf I}_{m})={\bf T}^{1/2}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})$. It follows from [51, rel. (14)] that

### v2

@@ VI-B Conditions of equality between efficient SFIM and parametric efficient FIM
By collecting the previous results, we are ready to state the following proposition whose proof is given in Appendix .9:
@@ VI-D The elliptical parameterized “low-rank” model
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\mathbf{0},\bm{\Sigma}_{0},g_{0})$ be a zero-mean, RES-distributed vector whose scatter matrix $\bm{\Sigma}_{0}$ is modeled as:
@@
  • $\mathbf{A}_{0}\triangleq\mathbf{A}(\bm{\gamma}_{0})\in\mathbb{R}^{m\times p}$ is a full rank column matrix with $m>p$, where $\bm{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{q}$ collects the parameter of interest and characterizes $\mathbf{A}_{0}$ and where the function $\mathbf{A}(\bm{\gamma})$ is differentiable,
@@ VII-A Real-complex representations
[…]
@@
    Use the complex Hilbert space (,,)($\mathcal{H}$$\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$) in (1), the “complex-aware” inner product $\left\langle h_{1},h_{2}\right\rangle_{\mathcal{H}}\triangleq{E_{0}\{h_{1}h_{2}^{*}\}}$ to express projection operators and tangent spaces,
@@
for NC-CES distributions where $\beta(g_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}\varphi^{2}_{c,0}(\mathcal{Q}_{c})\}}{m}$ with $\varphi_{c,0}(t)\triangleq\frac{-1}{g_{c,0}(t)}\frac{dg_{c,0}(t)}{dt}$ and $\mathbf{J}[\widetilde{\bm{\mu}}_{0}]]_{i,j}\triangleq\left.\frac{\partial[\widetilde{\bm{\mu}}(\bm{\theta})]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}}$. In the particular case of C-CES distributions, it is immediate to verify that the FIM in (128) reduces to
@@
${\bf J}[\mathrm{vec}({\bf A}_{0})]]_{i,j}\triangleq\left.\frac{\partial[\mathrm{vec}({\bf A}(\bm{\gamma}))]_{i}}{\partial\gamma_{j}}\right|_{\bm{\gamma}=\bm{\gamma}_{0}}$ and where here $\alpha(g_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}^{2}\varphi^{2}_{c,0}(\mathcal{Q}_{c})\}}{m(m+1)}$. Note that the parametric efficient FIM (131) reduces to:
@@ VIII Conclusion
[…]
    As a consequence, minimizing $\mathbf{h}-\mathbf{u}||_{\mathcal{H}_{q}}$ is equivalent to minimize each term $h_{i}-u_{i}||_{\mathcal{H}}$. Then, the equality (139) follows from the definition of orthogonal projection in $\mathcal{H}$ onto $\mathcal{U}$. ∎
@@
    where $\mathbf{G}(\mathbf{a},\mathbf{b})]_{i,j}\triangleq\left\langle a_{i},b_{j}\right\rangle_{\mathcal{H}}$.
@@
    To show that $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}$ is orthogonal to $p_{j}$, $\forall i,j$, we note that, according to its definition given in (18), $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}$ is the residual of $\mathbf{s}_{\bm{\gamma}_{0}}]_{i}$ after projection onto $\mathcal{T}_{2}+\mathcal{T}_{3}$. As a direct consequence, we have that $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}\perp(\mathcal{T}_{2}+\mathcal{T}_{3})$, or equivalently, $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}\in(\mathcal{T}_{2}+\mathcal{T}_{3})^{\perp}$. Moreover, again by definition in (25), $p_{j}\in(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}\subseteq\mathcal{T}_{2}+\mathcal{T}_{3}$. Consequently, $\left\langle[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i},p_{j}\right\rangle=0$, $\forall i,j$ since $(\mathcal{T}_{2}+\mathcal{T}_{3})\perp(\mathcal{T}_{2}+\mathcal{T}_{3})^{\perp}$.
@@
    By its definition, given in (13), $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$. As a consequence, we have that the first projection in the RHS of (148) can be evaluated as $\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}^{\perp})=\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$. And therefore $\mathbf{p}=\bm{0}\Leftrightarrow\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp})\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$ and $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{3}^{\perp}$. Then because $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$, $\mathbf{p}=\bm{0}\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{3}^{\perp}$ or equivalently, $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\perp\mathcal{T}_{3}$.
@@
    $\left(\frac{d[\mathbf{V}_{S}]_{11}}{d\underline{\mathrm{vecs}}(\mathbf{V}_{S})}\right)^{T}\triangleq\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}=-\frac{\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{[\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m}]_{1}}.$ (154)
@@
    $\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}=-\frac{\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{[\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m}]_{1}}.$ (156)
@@
    We note that, for $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$, we have $\mathbf{D}_{S}^{\bm{\Sigma}}=\mathbf{e}_{1,m}\mathbf{e}_{1,m}^{T}$, $\mathbf{D}_{S}^{\bm{\Sigma}}=m^{-1}\mathbf{I}_{m}$ and $\mathbf{D}_{S}^{\bm{\Sigma}}=m^{-1}|\bm{\Sigma}|^{1/m}\bm{\Sigma}^{-1}$ respectively.
@@
      $\mathbf{M}_{S}^{\bm{\Sigma}}]^{T}\underline{\mathrm{vecs}}(\mathbf{A})=\mathbf{D}_{m}\mathbf{K}_{\mathbf{V}_{S,0}}\underline{\mathrm{vecs}}(\mathbf{A})=\mathbf{D}_{m}\mathrm{vecs}(\mathbf{A})=\mathrm{vec}(\mathbf{A}).$ (158)
@@
      $\mathbf{M}_{S}^{\bm{\Sigma}}$ has full row rank equal to m(m+1)/21m(m+1)/2-1.
@@
    The proof of these properties can be found in [10, Lemma 4.2]. Let us now focus on the property P4 to add some insight on the image of $\mathbf{M}_{S}^{\bm{\Sigma}}$. From the Property P3, we have that the dimension of $\text{Im}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$, i.e. the number of the linearly independent rows of $\mathbf{M}_{S}^{\bm{\Sigma}}$ is equal to m(m+1)/21m(m+1)/2-1. Moreover, as a direct consequence of the rank-nullity theorem, we have the following additional property:
@@
      An orthonormal basis of the m(m+1)/21m(m+1)/2-1-dimensional image $\text{Im}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ of the restriction to $\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ of the linear application defined by $\mathbf{M}_{S}^{\bm{\Sigma}}$ is given by the columns of the matrix $\mathbf{V}_{\bm{\Sigma}}=\mathbf{D}_{m}\mathbf{U}_{\bm{\Sigma}}\in\mathbb{R}^{m^{2}\times(m(m+1)/2-1)}$ such that :
@@
    Derivation of the semiparametric nuisance tangent space $\mathcal{T}_{g_{0}}$ Let us consider the non-parametric model introduced in (47) as:
@@
    where $\mathbf{s}_{\bm{\rho}_{0},i}=\nabla_{\bm{\rho}}\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho}_{0},i})$ is the score vector of $\bm{\rho}_{0}\in\Upsilon_{i}$ in the ithi-th parametric sub-model $\mathcal{P}_{\tau_{i}}$. Consequently, according to [14] and [13, Sec. 4.4], the tangent space $\mathcal{T}_{g_{0}}$ can be defined as the closure  [^4] The closure $\overline{\mathcal{A}}$ of a set $\mathcal{A}$ is defined as the smallest closed set that contains $\mathcal{A}$, or equivalently, as the set of all elements in $\mathcal{A}$ together with all the limit points of $\mathcal{A}$. of the union of all the (parametric) tangent spaces $\mathcal{P}_{\tau_{i}}$:
@@
    Equivalently, $\mathcal{T}_{g_{0}}\subseteq\mathcal{H}$ is the subspace  [^5] The closure of a union of linear spaces doesn’t need to be linear, in general. However, as discussed in [48, Assumption S] and [13, Sec. 4.4, Remark 5]), $\mathcal{T}_{g_{0}}$ is a linear space in the vast majority of the non-pathological statistical models. of $\mathcal{H}$ composed by all the functions $h\in\mathcal{T}_{g_{0}}\subseteq\mathcal{H}$ for which there exists a sequence $\{\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i},\mathbf{c}_{i}\in\mathbb{R}^{r_{i}}\}_{i\in\mathbb{N}}$ such that $h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}}||^{2}=E_{0}\left\{(h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}})^{2}\right\}\rightarrow 0$.
@@
    Consequently, $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}$ is clearly an element of $\mathcal{T}_{g_{0}}$ thanks to the linearity of the expectation operator and to the fact that $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\}=0$ and $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]^{2}_{j}\}<+\infty$, for $j=1,\ldots,r_{i}$ from usual properties of the score vectors (under the regularity conditions discussed in [17, Sects. 6.2, 6.3]). Then $\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\in\mathcal{H}$ for $j=1,\ldots,r_{i}$ and it is measurable w.r.t. $Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})$, then $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}\in\mathcal{T}_{g_{0}},\forall i\in\mathbb{N}$.
@@
    $\begin{split}\int p_{X}(\mathbf{x}|\bm{\rho})d\mathbf{x}&=\int p_{0}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))d\mathbf{x}+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\int\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))p_{0}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))d\mathbf{x}\\ &=1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}{E_{0}\{\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\}}=1+0=1,\end{split}$ (170)
@@
    Consequently, since, as said before, any element of $\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$ is of the form $\mathbf{c}^{T}\mathbf{s}_{\bm{\rho}_{0},\tilde{i}}$ for some $\mathbf{c}\in\mathbb{R}^{r_{\tilde{i}}}$, we just need to choose $\mathbf{c}=\mathbf{e}_{j,r_{\tilde{i}}}$ to prove that $\mathcal{T}_{g_{0}}\ni\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\in\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$, where $\tilde{h}_{j}$ is a bounded function. The proof is completed by noticing that the set of bounded functions is dense in $\mathcal{H}$ and consequently, any element $h\in\mathcal{T}_{g_{0}}$ can be obtained as a converging sequence of bounded functions. This allow us to state that $\mathcal{T}_{g_{0}}\subset\overline{\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}}$ and conclude the proof. ∎
@@
    $\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]^{-1}=2\alpha^{-1}(g_{0})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{2\alpha^{-1}(g_{0})(\alpha(g_{0})-1)}{(m+2)\alpha(g_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}.$ (178)
@@
    Proof of Proposition VI.3. The different steps of the proof are based on some reasoning and notations of [42, Sec. VI of supplement material] for C-CES distributions and the general ”low-rank” model (131), that itself takes up the steps of the proof presented in [49].
@@
    which can be written, by means of the notation used in [42, Sec; V of supplement material], in the following compact form
@@
    with ${\bf u}_{n}={\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\mathrm{vec}({\bf I}_{m})={\bf T}^{1/2}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})$. It follows from [49, rel. (14)] that